OpenStudy (anonymous):

y=-2x+8 x=4 solve this, answer should be -2

OpenStudy (anonymous):

@ganeshie8 @HelpBlahBlahBlah @sleepyjess @sleepyhead314 @bohotness @YOLO_KING HELP please

OpenStudy (bohotness):

y=2x and y=-2x + 8 Both equations are in the format of... y=mx+b m is the slope. b is the y-intercept, the value of y when x=0. y=2x+0 slope is 2. y-intercept is 0, corresponding to the point of (0,0). Slope is change-of-y divided by change-of-x, also known as rise-over-run. For this equation, you will begin with the provided point and move the line such that it rises 2 units, while running 1 unit to the right. It will also descend 2 units, while running 1 unit to the left. y=-2x + 8 slope is -2. y-intercept is 8, (0,8). For this equation, you will begin with the provided point and move the line such that it rises 2 units, while running 1 unit to the left. It will also descend 2 units, while running 1 unit to the right. You can check your work by verifying that the two lines met at the correct coordinate... y=2x and y=-2x + 8 2x=-2x + 8 Let's add 2x to both sides of the equation... 2x+2x=2x-2x+8 4x=8 Divide both sides by 4... (4x)/4=8/4 x=2 Plug this into one equation to establish the value of y. The first equation is easiest... y=2x y=(2)(2) y=4 Plug both results into the other equation to verify... y=-2x + 8 4=[(-2)(2)]+8 4=-4+8 4=4 The two lines will meet at (2,4).

OpenStudy (sleepyjess):

Just substitute 4 for x

OpenStudy (sleepyjess):

Ummmmm..... what is all that @bohotness

OpenStudy (yolo_king):

Are we solving for X because I got no algebraic solution.

OpenStudy (sleepyjess):

Is it a system of equations? Or do you have to plug in the variable?

OpenStudy (anonymous):

look how i did it.....y=-2(4)+8=0 ,x=4 so ,0/4 answer is acc. to my solution

OpenStudy (bohotness):

okay

OpenStudy (anonymous):

but it's wrong....?

OpenStudy (sleepyjess):

Ok so you have to plug in the variable?

OpenStudy (anonymous):

here are 8 units rise and 4 units run and y=-2x+8 ...so how we will solve this @sleepyjess @bohotness

OpenStudy (bohotness):

y = 2x y = -2x + 8 y = 2x is a line through the origin with slope 2. y = -2x + 8 = -2(x - 4) is a line through (4,0) with slope -2. Since the lines have opposite slopes they, along with the x-axis, will form an isosceles triangle having a vertex at x = (0+4)/2 = 2. y = 2(2) = 4 y = -2(2)+8 = 4 Solution is (2,4). <_------- does that help

OpenStudy (sleepyjess):

@bohotness where are you getting y = 2x from???

OpenStudy (bohotness):

y=-2x+8 x=4 solve this, answer should be -2

OpenStudy (bohotness):

Y IS RIGHT THERE

OpenStudy (anonymous):

y's origion is ''0'' ,so rise is 8 and run is 4 if we make triangle then 8 to 4 rise over run so we will put the value of y =-2(4)+8=0 and x =4 then -2 is not the answer we getting???

OpenStudy (bohotness):

okay